OpenStudy (anonymous):

the length of a rectangle is three times its width. if the area of the rectangle 243 in^2, find its perimeter

OpenStudy (anonymous):

Area=Length*Width For this problem W=3L so, 234 in\[^{2}\] =L*3L \[243 \in ^{2}=3L ^{2}\] 243/3=L^2 L=9 The rest should be a standard perimeter equation relating the Length and Width that you just found.